OpenStudy (anonymous):

use a calculator and inverse trigonometric ratios to find the unknown measures.

OpenStudy (anonymous):

angle A angle B angle C AB BC AC

OpenStudy (anonymous):

(integral)ln(x) dx set u = ln(x), dv = dx then we find du = (1/x) dx, v = x substitute (integral) ln(x) dx = (integral) u dv and use integration by parts = uv - (integral) v du substitute u=ln(x), v=x, and du=(1/x)dx = ln(x) x - (integral) x (1/x) dx = ln(x) x - (integral) dx = ln(x) x - x + C = x ln(x) - x + C.